# DPImageBench: A Unified Benchmark for Differentially Private Image Synthesis

### v2

@@ 7.2 Model Architecture and Privacy Budget (RQ2)
Experiment Design. We study the utility and fidelity of synthetic images of studied methods across five distinct privacy budgets . This RQ also studies how the DP image synthesis methods perform under different sizes of synthesizers. We further evaluate the performance of DP image synthesis methods with varying synthesizer sizes under a certain privacy budget . For diffusion-based synthesizers, we consider sizes of , while for the model size of GAN-based synthesizers, we use .
Result Analysis. Figure 5 presents the utility and fidelity evaluations of synthetic images on the FashionMNIST dataset under different privacy budgets. Figure 6 presents the evaluations of synthetic images for CIFAR-10 across different model sizes. The results, presented in Fig 5 and Figure 6, highlight the following key takeaways:
@@
Adding noise on low-dimensional features is less sensitive to privacy budgets. Figure 5 shows that adding noise to high-dimensional features, like weight gradients, generally yields superior synthesis performance compared to adding noise to low-dimensional features, such as adding noise to training loss in DP-Kernel, especially when the privacy budget is large (e.g., and ). However, methods that add noise to high-dimensional features are more sensitive to the privacy budget. At very low privacy budgets, such as , methods that add noise to low-dimensional features outperform, with DP-Kernel and DP-MERF achieving the highest accuracy of 71.2% and 70.6%. These low-dimensional features exhibit lower sensitivity, allowing for a smaller noise scale than the noise scale added to high-dimensional features.
Batch size | M | M | M | M |
4,096 | 78.4 | 78.2 | 76.0 | 79.5 |
16,384 | 78.5 | 78.6 | 78.8 | 79.5 |
@@
Larger synthesizer models do not always yield better performance. GANs often encounter unstable training [39, 83] and mode collapse issues [84, 85]. Training a larger GAN model under DP is even more challenging. As a result, as seen in Figure 6, scaling up model sizes does not necessarily benefit GAN-based methods. A similar phenomenon can be observed in diffusion-based methods. For instance, in PrivImage, the accuracy of synthetic images decreases from 78.2% to 76.0% as the synthesizer size grows from 11.1M to 19.6M. Although larger synthesizers can better learn to generate high-quality synthetic images, the noise added to the gradient increases with model size under the same privacy budget, which can potentially degrade synthetic performance. Li et al. [4] demonstrated that when a synthesizer reaches the upper limit of its generalization capability—where increasing the model’s parameter size no longer improves synthesis performance—DP can further degrade results, with larger synthesizers potentially performing worse. One solution is to increase the batch size of DP-SGD to reduce the DP noise scale [86]. Table XI presents the Acc of synthetic images from PrivImage with batch sizes of 4,096 and 16,384 (the batch size used in PrivImage [4]). We observe that the large batch size of DP-SGD is better for large model sizes.
7.3 Leveraging Public Image Datasets (RQ3)
@@
The default pretraining strategy is introduced in Appendix B.2.3. As described in Section 4.1.2, GS-WGAN pretrains synthesizers using the sensitive dataset without the need for external public datasets, unlike other methods such as PrivImage [4], PDP-Diffusion [32], and DP-LDM [18]. PrivImage and PDP-Diffusion both build upon DPDM but use a distinct pretraining method. Without pretraining, they both revert to the original DPDM.
PDP-Diffusion | 25K | 50K | 100K | 200K | 400K |
Acc (%) | 97.4 | 97.6 | 97.4 | 97.4 | 97.1 |
FLD | 2.48 | 2.56 | 2.83 | 3.40 | 3.00 |
PrivImage | 25K | 50K | 100K | 200K | 400K |
Acc (%) | 97.8 | 97.4 | 97.5 | 97.8 | 97.2 |
FLD | 2.23 | 2.48 | 2.62 | 2.80 | 2.90 |
Result Analysis. The key takeaways from this section are as follows:
@@
FLD | 5% | 5% (rand.) | 5%+5% (rand.) | 10% (rand.) | Whole |
DP-MERF | 30.2 | 31.6 | 29.8 | 29.7 | 28.4 |
DP-NTK | 50.0 | 50.5 | 55.9 | 53.6 | 49.8 |
DP-Kernel | 31.5 | 30.1 | 30.3 | 31.1 | 30.0 |
DP-GAN | 18.9 | 18.9 | 21.8 | 21.6 | 29.5 |
PE | 4.5 | 6.5 | 12.2 | 10.6 | 8.5 |
DP-FETA | 10.7 | 12.4 | 9.3 | 12.8 | 10.6 |
PDP-Diffusion | 4.7 | 10.7 | 6.1 | 9.3 | 7.2 |
DP-LDM (SD) | 5.1 | 11.7 | 7.2 | 10.6 | 9.0 |
DP-LDM | 12.2 | 15.2 | 13.0 | 14.4 | 14.1 |
DP-LoRA | 8.1 | 11.5 | 8.0 | 11.4 | 9.3 |
Pretraining is more beneficial for diffusion-based methods than for GAN-based methods. Figure 7 shows the Acc and FLD of synthetic images for CIFAR-10, comparing results with and without pretraining, and examining both conditional and unconditional pretraining methods. The bar on the left indicates lower values, while the bar on the right indicates higher values. This figure shows that for diffusion-based methods, pretraining consistently improves both the utility and fidelity of synthetic images, while its impact on GAN-based methods varies. For example, for DP-Kernel, the Acc shows an 8.2% reduction while pretraining the synthesizer. GAN-based models have limited ability to produce high-quality images, so pretraining has less effect on their performance than it does for diffusion-based methods.
Conditional pretraining significantly outperforms unconditional pretraining. For example, in Figure 7, using unconditional pretraining for PrivImage results in a 34.4% drop in accuracy and a 2.1 increase in FLD. The reason is that using labels for conditional pretraining helps establish a stronger connection between the labels and the images. On the other hand, when using the entire public dataset for pretraining, the presence of various labels unrelated to the sensitive images can cause the model to learn redundant label-to-image mappings. This distraction can compromise its ability to learn the meaningful and relevant features necessary for accurately synthesizing sensitive images. Table XII also shows that excessive pretraining may reduce the performance of synthetic images.
A closer distribution match between the pretraining distribution and the sensitive dataset improves DP image synthesis. PrivImage, which selects a subset of public images whose distribution closely matches that of the sensitive images, produces synthetic images with the highest utility on CIFAR-10 in Table VIII. To further investigate the impact of the relationship between pretraining datasets and sensitive datasets on DP synthesis performance, Figure 8 compares the Acc and FLD for the sensitive datasets FashionMNIST and CIFAR-10 when pretraining with public datasets ImageNet and Places365. The average accuracy scores for CIFAR-10 using ImageNet and Places365 as the pretraining datasets are 48.3 and 42.3, respectively. For FashionMNIST, the average accuracy scores with ImageNet and Places365 are 77.2 and 76.1. Actually, ImageNet’s distribution is closer to CIFAR-10 than Places365 because its object-focused categories align more with CIFAR-10’s content [59]. Places365, focused on scenes [68], introduces less relevant background information, making its distribution less similar to CIFAR-10’s object-centric data. For GAN-based methods, pretraining offers minimal improvement, with performance varying slightly depending on whether ImageNet or Places365 is used. In contrast, for diffusion-based methods, pretraining enhances performance, as shown in Figure 7. Pretraining on ImageNet consistently outperforms Places365, except in the case of PDP-Diffusion trained on FashionMNIST, where the difference is marginal (0.3%).
7.4 Combining Algorithmic Improvement (RQ4)
Experiment Design. This RQ explores the potential of integrating the strengths of various methods to enhance existing approaches, laying the groundwork for future DP image synthesis methods.
Combining PrivImage with others. We construct four subsets from the public dataset ImageNet and compare the synthesis performance when using different subsets for pretraining. These subsets are obtained by (1) randomly selecting 5% of images; (2) randomly selecting 10% of images; (3) using the distribution matching method in PrivImage to select 5%, ensuring that the chosen subset matches the semantic distribution of the sensitive dataset; and (4) combining 5% from PrivImage with an additional 5% randomly selected. PrivImage is equivalent to PDP-Diffusion with 5% selected pretraining data.
Combining DP-FETA with others. We incorporate a shortcut process into the DP training, extending it to other methods. Specifically, we extract central images from sensitive datasets and train synthesizers on these images prior to training on the sensitive images. Central images are perturbed with Gaussian noise for DP. PE is a fine-tuning-free method, so we exclude it from this analysis.
Privacy Analysis. Since both the data selection of PrivImage and central image constructions of DP-FETA will introduce additional privacy costs, combining them with other methods needs privacy composition. Considering the nice linearly composability property of Renyi DP (RDP) [34], we implement other methods with RDP, and use RDP to compose the privacy costs of two processes.
Result Analysis. Table XIII presents the results of using just 5% of a public dataset, aligned with the categories of CIFAR-10. Table XIV shows the results of integrating the DP training shortcut into other methods. These results highlight the following key takeaways:
Integrating the pretraining dataset selection proposed in PrivImage can enhance synthetic performance. The pretraining dataset selection improves the similarity between the public dataset and the sensitive dataset, and leads to accuracy improvements of 8.8% () for PDP-Diffusion and 4.9% () for DP-LDM (SD), respectively. For diffusion-based methods, the FLD of synthetic images is consistently better when pretraining synthesizers with 5% of a public dataset compared to using the entire dataset. Compared to randomly selecting images from the public dataset, there are performance decreases in Acc and FID when using the entire public dataset for pretraining. The improvement for GAN-based methods is minimal because GANs struggle more than DMs to learn the distribution of the public dataset. These results indicate that the relationship between the distributions of public and sensitive datasets plays a crucial role in improving synthetic image performance, which aligns with findings in other data domains [13]. This experiments also present efforts that incorporate the strength of selecting public datasets as proposed taxonomy in Section 4 to enhance methods.
Building DP training shortcut may damage the synthetic performance. Table XIV shows that incorporating a DP training shortcut may impair the synthetic performance of methods with pretraining. For instance, the Acc and FLD of DP-LDM (SD) decline from 81.6% and 11.7 to 80.6% and 13.3 when shortcuts are added. This is because the shortcut includes training on the central images after pretraining, which may damage the pretraining synthesizer. For methods without pretraining, the fidelity of synthetic images (measured by FLD) is consistently improved, while the improvement of utility (measured by Acc) fluctuates.
@@ 7.4 Combining Algorithmic Improvement (RQ4)
We hypothesize that the DP shortcut may disrupt the pretraining process, whereas pretraining dataset selection does not affect other algorithmic components. We need further research to effectively integrate the strengths of diverse algorithms.
Algorithm | With DP Training Shortcuts | Original | ||
(CIFAR-10) | Acc (%) | FLD | Acc (%) | FLD |
DP-MERF | 57.7 | 24.5 | 62.2 | 29.2 |
DP-NTK | 70.6 | 31.0 | 76.3 | 36.4 |
DP-Kernel | 78.0 | 18.5 | 70.0 | 21.3 |
GS-WGAN | 58.9 | 27.2 | 56.7 | 28.1 |
DP-GAN | 55.4 | 22.8 | 70.3 | 23.9 |
DPDM | 87.3 | 4.6 | 85.6 | 6.6 |
PDP-Diffusion | 85.0 | 4.7 | 85.4 | 4.9 |
DP-LDM (SD) | 80.6 | 13.3 | 81.6 | 11.7 |
DP-LDM | 85.3 | 18.8 | 86.3 | 15.4 |
DP-LoRA | 86.3 | 14.2 | 83.8 | 14.8 |
PrivImage | 86.4 | 4.6 | 87.1 | 4.3 |
Average | 75.6 | 16.7 | 76.8 | 17.9 |
8 Related Work
@@
Several studies have established benchmarks for studying the role of DP in preserving the privacy of machine learning models [3, 87, 1, 21, 22]. However, Du et al. [3] and Tao et al. [87] focus exclusively on DP tabular data synthesis. Hu et al. [1] publish a systematization of knowledge on DP dataset synthesis covering various types of data, including tabular and graph data, among others. However, they overlook DP image synthesizers involving pretraining [4, 24] and foundational APIs [5]. Ganev et al. [22] propose a benchmark tailored for PATE-GAN. DPImageBench is the first benchmark to incorporate state-of-the-art synthesizers, such as DMs, and pretraining techniques for the DP image synthesis field.
9 Conclusions
@@
DP image synthesis faces challenges due to complex codebases, inconsistent hyper-parameters, diverse model architectures, and the absence of standardized evaluation metrics. Advancing DP image synthesis algorithms requires a thorough understanding of existing methods and fair, direct comparisons. This paper introduces DPImageBench, a benchmark for evaluating DP image synthesis algorithms. First, in terms of methods, we analyze and systematically characterize twelve prominent approaches, detailing their model architectures, pretraining strategies, and privacy mechanisms. Our novel taxonomy examines where the noise is added to satisfy DP, enabling the exploration of combined DP image synthesis methods. Second, for evaluation, we use seven private datasets, two public datasets, and seven fidelity and utility metrics. Our findings reveal that selecting downstream classifiers based on the highest accuracy on the sensitive test set violates DP and inflates utility scores. DPImageBench addresses and corrects these issues. Third, the platform offers a standardized interface, enabling unified implementation for current and future methods.
Our analysis yields several key insights. Notably, we highlight that pretraining on public image datasets does not universally improve performance. Instead, the distributional similarity between pretraining and sensitive images plays a crucial role and can sometimes limit the benefits. Besides, we find that current methods still face challenges in high-resolution DP image synthesis. These findings highlight potential research opportunities for optimizing pretraining strategies to enhance DP image synthesis.
@@ Appendix A Privacy Notations
As shown in Table I, different definitions of neighboring datasets (i.e., replace-one, add-or-remove-one) have been used in prior work. In this section, we show that the methods that used the replace-one definition—DP-MERF, DP-NTK, and DP-Kernel—can all be adapted to the add-or-remove-one definition without affecting their DP parameters.
Our key idea is to show that the sensitivity of the function these methods privatize remains the same when switching from replace-one to add-or-remove-one. For all these three methods, the function to be privatized is the mean (vector or kernel) feature embedding of the private dataset. For simplicity, we will use DP-MERF as an example, and the proof for the other two methods is similar. Let be the number of private samples, and be the feature of sample . Let the two neighboring datasets (under the replace-one definition) be and where they only differ on the -th sample: . Eq. (9) and (10) in DP-MERF [28] show that the sensitivity of the mean feature embedding is :

### v3

@@ 7.2. Model Size and Privacy Budget (RQ2)
Experiment Design. We evaluate the utility and fidelity of synthetic images across six privacy budgets . We further evaluate the performance of DP image synthesis methods with varying synthesizer sizes under a certain privacy budget . For diffusion-based synthesizers, we consider sizes of , while for the model size of GAN-based synthesizers, we use .
Result Analysis. We summarize the key takeaways from this section as follows:
@@
Adding noise on low-dimensional features is less sensitive to privacy budgets. Figure 5 shows that adding noise to high-dimensional features, like weight gradients, generally yields superior synthesis performance compared to adding noise to low-dimensional features, such as adding noise to training loss in DP-Kernel, especially at large privacy budgets like and . However, methods that add noise to high-dimensional features are more sensitive to the privacy budget. At very low privacy budgets, such as , methods that add noise to low-dimensional features outperform, with DP-Kernel and DP-MERF achieving the highest accuracy of 71.2% and 70.6%, due to their lower sensitivity and smaller noise scale.
@@
Larger synthesizer models do not always yield better performance. GANs often face unstable training (Arjovsky et al., 2017; Lin et al., 2021b) and mode collapse (Lin et al., 2018; Goodfellow, 2016), with DP further complicating training of larger models. Figure 4 shows that scaling up model sizes does not always benefit GAN-based methods. Similarly, in diffusion-based methods like PrivImage, accuracy drops from 78.2% to 76.0% as synthesizer size increases from 11.1M to 19.6M. Larger synthesizers, despite better learning capacity, suffer from increased DP noise under the same privacy budget, potentially degrading performance. Li et al. (Li et al., 2024) noted that when synthesizers reach their generalization limit, DP exacerbates performance decline in larger models. Increasing the batch size in DP-SGD can mitigate DP noise (De et al., 2022), as shown in Table 19 (Appendix D.4), where PrivImage achieves better Acc with a batch size of 16,384 (used in PrivImage (Li et al., 2024)) compared to 4,096, especially for large model sizes.
7.3. Leveraging Public Image Datasets (RQ3)
@@
Appendix B.2.3 introduces the default pretraining strategy. As noted in Section 4.1.2, GS-WGAN pretrains synthesizers using sensitive datasets, unlike PrivImage (Li et al., 2024), which rely on external public datasets. PrivImage and PDP-Diffusion, built on DPDM, use a distinct pretraining method, reverting to DPDM without pretraining.
Result Analysis. The takeaways from this section are as follows:
Pretraining is more beneficial for diffusion-based methods than for GAN-based methods. Figure 6 compares the Acc and FLD of synthetic CIFAR-10 images with and without pretraining, using both conditional and unconditional methods. Lower values are shown on the left bar and higher on the right. It reveals that pretraining consistently enhances utility and fidelity in diffusion-based methods, but its effect on GAN-based methods varies. For DP-Kernel, the Acc shows an 8.2% reduction while pretraining the synthesizer. GAN-based models have limited ability to produce high-quality images, so pretraining has less effect on their performance than it does for diffusion-based methods.
Conditional pretraining outperforms unconditional pretraining. In Figure 6, using unconditional pretraining for PrivImage results in a 34.4% drop in accuracy and a 2.1 increase in FLD. The reason is that using labels for conditional pretraining helps establish a stronger connection between the labels and images. When using entire public datasets for pretraining, the presence of various labels unrelated to the sensitive images can cause the model to learn redundant label-to-image mappings. This distraction compromises its ability to learn the meaningful features necessary for accurately synthesizing sensitive images. Table 20 of Appendix D.5 shows that excessive pretraining may reduce the synthetic performance.
A closer distribution match between pretraining and sensitive datasets enhances DP image synthesis. PrivImage, which uses a public image subset closely matching the sensitive images’ distribution, achieves the highest utility on CIFAR-10 in Table 6. Figure 7 compares the Acc and FLD for the sensitive datasets CIFAR-10 when pretraining with public datasets ImageNet and Places365. Appendix D.5 of Figure 11 shows the results of FashionMNIST. The average accuracy scores for CIFAR-10 using ImageNet and Places365 as the pretraining datasets are 48.3 and 42.3, respectively. Actually, ImageNet’s distribution is closer to CIFAR-10 than Places365 because its object-focused categories align more with CIFAR-10’s content (Deng et al., 2009). Places365, focused on scenes (Zhou et al., 2017), introduces less relevant background information, making its distribution less similar to CIFAR-10’s object-centric data. For GAN-based methods, pretraining offers minimal improvement, with performance varying slightly depending on whether ImageNet or Places365 is used. In contrast, for diffusion-based methods, pretraining enhances performance, as shown in Figure 6.
7.4. Combining Algorithmic Improvement (RQ4)
Experiment Design. We create four subsets from ImageNet to compare synthesis performance during pretraining: (1) 5% randomly selected images; (2) 10% randomly selected images; (3) 5% selected via PrivImage’s distribution matching to align with the sensitive dataset’s semantics; and (4) 5% from PrivImage plus an additional 5% randomly selected. PrivImage is equivalent to PDP-Diffusion with 5% selected pretraining data.
Result Analysis. Table 9 presents the results of using just 5% of a public dataset, aligned with the categories of CIFAR-10. Table 22 in Appendix shows the results of FLD. This table highlights the following key takeaway:
Integrating the pretraining dataset selection proposed in PrivImage can enhance synthetic performance. The pretraining dataset selection improves the similarity between the public and sensitive dataset, boosting accuracy by 8.8% (78.9%–70.1%) for PDP-Diffusion and 4.9% (74.8%–69.9%) for DP-LDM (SD). Compared to randomly selecting images from the public dataset, there are performance decreases in Acc when using the entire public dataset for pretraining. The improvement for GAN-based methods is minimal because GANs struggle more than DMs to learn the distribution of the public dataset. These results indicate that the relationship between the distributions of public and sensitive datasets plays a crucial role in improving synthetic image performance, which aligns with findings in other data domains (Yin et al., 2022). This experiments also present efforts that incorporate the strength of selecting public datasets as proposed taxonomy in Section 4 to enhance methods.
@@ 7.4. Combining Algorithmic Improvement (RQ4)
Appendix D.6 presents the results of combining the DP shortcut proposed in DP-FETA (Li et al., 2025) with other methods, showing that algorithmic combinations do not always enhance synthetic performance. We hypothesize that the DP shortcut may disrupt the pretraining process, whereas pretraining dataset selection does not affect other algorithmic components. We need further research to effectively integrate the strengths of diverse algorithms.
Privacy Analysis. Since both the data selection of PrivImage and central image constructions of DP-FETA will introduce additional privacy costs. We use Renyi DP (Mironov, 2017) to compose privacy costs.
8. Related Work
@@
Several studies provide benchmarks for DP in machine learning (Tao et al., 2021; Hu et al., 2024; Wei et al., 2023; Ganev et al., 2024). Du et al. (Du and Li, 2024) and Tao et al. (Tao et al., 2021) focus on DP tabular data synthesis, while Hu et al. (Hu et al., 2024) cover various data types (e.g., tabular, graph) but overlook DP image synthesizers with pretraining (Li et al., 2024; Wang et al., 2024) and foundational APIs (Lin et al., 2024). DPImageBench is the first benchmark to include state-of-the-art synthesizers (e.g., DMs and APIs) and pretraining for DP image synthesis.
9. Conclusions
@@
DP image synthesis faces challenges from complex codebases, inconsistent hyper-parameters, diverse architectures, and lacking standardized evaluation metrics. Advancing these algorithms requires understanding existing methods and fair comparisons. This paper introduces DPImageBench, a benchmark for DP image synthesis evaluation. First, for methods, we analyze twelve prominent approaches, detailing their architectures, pretraining, and privacy mechanisms, using a novel taxonomy to explore noise addition for DP and combined methods. Second, for evaluation, we use seven private and two public datasets with seven fidelity and utility metrics, revealing that selecting classifiers based on the highest accuracy on sensitive test sets violates DP and inflates utility scores—issues DPImageBench addresses. Third, the platform provides a standardized interface for the unified implementation of current and future methods. Our analysis yields several key insights. The pretraining on public image datasets does not always improve performance; the distributional similarity between pretraining and sensitive images is key and can limit benefits. Besides, current methods still face challenges in high-resolution DP image synthesis. These findings highlight potential research for optimizing pretraining strategies.
@@ Appendix A Privacy Notations
As shown in Table 12, different definitions of neighboring datasets (i.e., replace-one, add-or-remove-one) have been used in prior work. In this section, we show that the methods that used the replace-one definition—DP-MERF, DP-NTK, and DP-Kernel—can all be adapted to the add-or-remove-one definition without affecting their DP parameters.
Our key idea is to show that the sensitivity of the function these methods privatize remains the same when switching from replace-one to add-or-remove-one while the dataset size is unknown. For all these three methods, the function to be privatized is the mean (vector or kernel) feature embedding of the private dataset. For simplicity, we will use DP-MERF as an example, and the proof for the other two methods is similar. Let be the number of private samples, and be the feature of sample . Let the two neighboring datasets (under the replace-one definition) be and where they only differ on the -th sample: . Eq. (9) and (10) in DP-MERF (Harder et al., 2021) show that the sensitivity of the mean feature embedding is :
| DP-MERF | 85.0 | 82.5 | 63.7 | 61.8 | 22.8 | 25.3 | 1.8 | 1.8 | 27.8 | 39.2 | 79.2 | 76.3 | 67.9 | 67.1 |
| DP-NTK | 62.0 | 84.0 | 59.7 | 75.4 | 17.2 | 23.5 | 1.3 | 4.2 | 25.4 | 31.6 | 61.8 | 61.7 | 55.0 | 51.5 |
| DP-Kernel | 95.7 | 94.3 | 76.7 | 78.2 | 30.4 | 28.0 | 6.7 | 6.9 | 51.0 | 50.0 | 81.2 | 84.9 | 69.8 | 72.1 |
| Algorithm | MNIST | FashionMNIST | CIFAR-10 | |||||||||||||||
| DP-MERF | 72.8 | 2.35 | 0.03 | 0.23 | 27.4 | -2.25 | 78.8 | 3.04 | 0.15 | 0.00 | 23.9 | -1.94 | 181.8 | 3.41 | 0.26 | 0.00 | 27.5 | -2.28 |
| DP-NTK | 31.4 | 2.16 | 0.25 | 0.34 | 15.8 | -2.14 | 90.0 | 3.41 | 0.05 | 0.03 | 26.9 | -2.02 | 142.0 | 3.33 | 0.32 | 0.00 | 25.9 | -2.28 |
| DP-Kernel | 43.9 | 2.22 | 0.11 | 0.01 | 22.4 | -2.15 | 64.2 | 2.82 | 0.19 | 0.00 | 17.3 | -1.93 | 201.4 | 2.90 | 0.31 | 0.00 | 30.4 | -2.28 |
| Algorithm | CIFAR-100 | EuroSAT | CelebA | |||||||||||||||
| DP-MERF | 185.6 | 3.02 | 0.36 | 0.00 | 24.6 | -2.28 | 211.9 | 2.14 | 0.12 | 0.00 | 34.8 | -2.27 | 129.0 | 2.76 | 0.13 | 0.00 | 15.0 | -2.01 |
| DP-NTK | 440.2 | 1.26 | 0.00 | 0.00 | 45.4 | -2.28 | 156.6 | 2.57 | 0.10 | 0.00 | 29.8 | -2.21 | 200.0 | 3.01 | 0.05 | 0.00 | 33.7 | -2.00 |
| DP-Kernel | 159.2 | 2.95 | 0.34 | 0.00 | 23.0 | -2.28 | 163.4 | 2.42 | 0.11 | 0.00 | 30.1 | -2.25 | 113.9 | 2.45 | 0.13 | 0.00 | 14.4 | -1.83 |
@@
Previous works (Abadi et al., 2016; Kulesza et al., 2024) assume a known dataset size and overlook the privacy budget required for dataset size estimation. Similarly, DPImageBench methods using DP-SGD, including PrivImage (Li et al., 2024), DP-LoRA (Tsai et al., 2024), and DP-FETA (Li et al., 2025), also assume is known. Then, the sensitivity of the function is defined as follows,
This equation shows that the sensitivity is halved when the dataset size is known. We adjust the sensitivity for DP-MERF, DP-NTK, and DP-Kernel, and present the results in Table 10 and Table 11. The reduced sensitivity can reduce the noise scale under a certain privacy budget. 14 out of 18 entries in Table 7 surpass the results in Table 6 in terms of FID. In Table 6, we observe that the improvement in Acc is not consistent when compared to the results in Table 6 (27 out of 42 entries in Table 10 exceed the results in Table 6). This inconsistency arises because the quality of the synthetic images is not sufficiently high, leading to instability in the Acc.
| Algorithm | Privacy notion |
| DP-MERF (Harder et al., 2021) | Replace-one |
| DP-NTK (Yang et al., 2023) | Replace-one |
| DP-Kernel (Jiang et al., 2023) | Replace-one |
| PE (Lin et al., 2024) | Add-or-remove-one |
| GS-WGAN (Chen et al., 2020) | Both |
| DP-GAN (Xie et al., 2018) | Add-or-remove-one |
| DPDM (Dockhorn et al., 2023) | Add-or-remove-one |
| DP-FETA (Li et al., 2025) | Add-or-remove-one |
| PDP-Diffusion (Ghalebikesabi et al., 2023) | Add-or-remove-one |
| DP-LDM (Liu et al., 2023) | Add-or-remove-one |
| DP-LoRA (Tsai et al., 2024) | Add-or-remove-one |
| PrivImage (Li et al., 2024) | Add-or-remove-one |